OpenStudy (anonymous):

I'm trying to find what the summation from 0 to infinity of ((e^-4)4^x)/4! converges to (trying to show it is a valid pdf). Any suggestions on how to approach this would be welcome.

OpenStudy (jhannybean):

\[\huge \sum_{0}^{\infty} \frac{4^x(e^{-4})}{4!}\] Is it this? haha

OpenStudy (anonymous):

haha yes

OpenStudy (jdoe0001):

hmmm dunno that one :(

OpenStudy (amistre64):

does the sequence limit to zero?

OpenStudy (amistre64):

e^(-4)/4! is just some constant value ...

OpenStudy (anonymous):

I believe it does, since the factorial grows faster than the exponential

OpenStudy (anonymous):

oh sorry it's over x!

OpenStudy (amistre64):

4^x is the only thing that varies .. the constant factors out ...

OpenStudy (anonymous):

you're right, without that it wouldn't converge

OpenStudy (amistre64):

.... can you pic the question? avoids alot of error

OpenStudy (anonymous):

will try, h/o

OpenStudy (anonymous):

\[p(x) = \frac{e^{-4}4^x}{x!}\]

OpenStudy (amistre64):

pdf, probability density function ...

OpenStudy (anonymous):

yeah, pmf not pdf. sorry, the pic was coming out really grainy

OpenStudy (anonymous):

\[x = 0,1,2,...\]

OpenStudy (amistre64):

we want the sumto equal ...1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i know i can pull out the constant

OpenStudy (amistre64):

then it behooves us to consider that 4^x/x! = e^(4)

OpenStudy (amistre64):

sum of ...

OpenStudy (amistre64):

do you recall the taylor series for e^x ?

OpenStudy (anonymous):

i don't i can look it up, that may be the prod in the right direction i am looking for

OpenStudy (amistre64):

its a thought, not sure how well it will turn out tho

OpenStudy (anonymous):

ah, yes. that looks like it exactly

OpenStudy (anonymous):

\[e^x = \sum \frac{x^n}{n!}\]

OpenStudy (amistre64):

im so clever it scares me ....

OpenStudy (anonymous):

haha i liked how you looked at it in terms of solving for what would give us e^4

OpenStudy (anonymous):

i appreciate the help

OpenStudy (amistre64):

youre welcome :)